please these questions are reakted to the graphs below. answer the following questions.

-calculate the twitch force to tetanic force ratio?
-what is the frequency at which the first indication of force summation occurs
-what is the frequency at which a fused tetanic force response is recorded?

Muscle tension vs. Voltage Muscle tension (g) 0.05 0.1 0.2 0.25 0.3 0.15 Voltage (Volts)
Muscle tension vs Frequency Muscle tension (g) 1 10 20 40 50 60 30 Frequency (Hz)

related*
Answer #1

1.

Table 1:

Let points in the graph as g1, g2, ......., g6

Table 2:

Let the points in the graph as m1, m2, ......, m5

Twitch Force = (g1 - g2 ) / (0.2- 0.1)

= (18 - 7 ) / 0.1

= 11 / 0.1

= 110

Tetanic Force = (m5 - m4 ) / (50-10)

= (170 - 100) / 40

= 70 / 40

= 1.75

Twitch force / Tetanic force = 110 / 1.75

= 62.85

2. First indication of force summation occurs from 0.1 to 0.2 Volts because in force summation the muscle tension increases with every stimulus.

3. Fequency at which fused tetanic force occurs at 20 Hz because muscle contraction sustains even at 50Hz.